OpenStudy (raiders88):

find the simplified form of the expression. give your answer in scientific notation (5 x 10^-7) (8 x 10^-4)

OpenStudy (daniellelovee):

do you have a guess?

OpenStudy (misty1212):

no, no guess

OpenStudy (daniellelovee):

lol yeah he/she went offline

OpenStudy (raiders88):

hey so i need help with more

OpenStudy (jango_in_dtown):

(5*10^-7)(8*10^-4)=(5*8)*10^-11=40*10^-11=4*10^(-10)

OpenStudy (jango_in_dtown):

@raiders88

OpenStudy (raiders88):

no need to be on here its closed